# Self-supervised Pre-training for Mapping of Archaeological Stone Wall in Historic Landscapes Using High-Resolution DEM Derivatives

### v1

@@ Abstract
Dry-stone walls hold significant heritage and environmental value. Mapping and maintaining these structures is essential for preserving ecosystems and supporting wildfire control during dry seasons in Australia. However, a substantial portion of stone walls located in wild landscapes remain unidentified and unregistered due to their inaccessibility and the high cost of mapping. Deep learning-based segmentation models offer a scalable solution for automated mapping of dry-stone walls, but two challenges remain, including (1) visual occlusion of low-lying stone walls by dense vegetation and (2) limited availability of labeled data for training supervised segmentation models. In this study, we propose a segmentation approach for the automatic and accurate mapping of low-lying dry-stone walls using high-resolution digital elevation models (DEMs). The use of Airborne LiDAR-derived DEMs help overcome the visual occlusion problem caused by dense vegetation and canopy cover, by capturing underlying terrain structures that are invisible in traditional optical imagery. Our approach proposes DINO-CV, a self-supervised cross-view pre-training strategy based on knowledge distillation, to address the challenge of limited labeled data. DINO-CV learns invariant visual and structural features across multiple high-resolution DEM derivatives, enabling accurate stone wall mapping on various computer vision backbones (i.e., ResNet, Wide ResNet, and Vision Transformers (ViTs)). Applied to the UNESCO World Heritage cultural landscape of Budj Bim, Victoria, our method successfully identifies one of Australia’s densest collections of colonial dry-stone walls beyond Indigenous cultural heritage. The proposed method achieves a mean Intersection over Union (mIoU) of 68.6% between predicted and ground-truth stone wall locations in the test areas, and maintains robust performance of 63.8% mIoU when fine-tuned with limited labeled data (10%) using self-supervised vision backbones pre-trained on the DEM derivatives of Budj Bim landscape. These results demonstrate the potential of applying self-supervised learning on high-resolution DEM derivatives for automated mapping of dry-stone walls, particularly in densely vegetated environments and heritage landscapes with limited labeled data.
@@
Mapping of historic dry-stone walls , Self-supervised learning , DEMs segmentation , Wildfire control
@@ 1 Introduction
Traditional methods of mapping dry-stone walls, such as manual surveys or aerial photography, can be impractical and labor-intensive, especially in remote or rugged terrain. Recent advancements in computer vision and deep learning have enabled the development of automated, scalable mapping techniques that can efficiently detect dry-stone walls from high-resolution imagery and LiDAR data (Trotter et al., 2022; Suh and Ouimet, 2023; Wang et al., 2024). Supervised deep neural networks, such as ConvNets (CNNs) (He et al., 2016), vision transformers (ViTs) (Dosovitskiy et al., 2021), and U-Net architectures (Ronneberger et al., 2015), can be trained on annotated datasets to learn the visual and spatial features of dry-stone walls, including their shape, texture, and spatial patterns. These deep-learning based mapping approaches largely reduce the time and cost of mapping and offer the potential for large-scale registration of undocumented structures. However, the performance of these deep supervised models heavily depends on the availability of large, labeled datasets for model training, which are often difficult and costly to obtain. Another critical challenge is the visual occlusion of low-lying stone wall structures by dense vegetation in wild landscapes, which makes it difficult for models to accurately detect and localize these features.
@@
Grounded in these hypotheses, this study poses the following research question: How can self-supervised learning on high-resolution DEM derivatives improve the mapping performance of low-lying dry-stone walls in complex, vegetated environments? To answer this question, we explore the potential of self-supervised learning on high-resolution DEM derivatives for automated mapping of dry-stone walls in complex, vegetated environments. Hence, we propose a self-supervised cross-view pre-training strategy, DINO-CV, based on knowledge distillation to learn visual and structural features from multiple DEM derivatives generated from high-resolution aerial LiDAR data. The mapping performance of our approach is evaluated on the Budj Bim cultural landscape in Victoria, Australia, a UNESCO World Heritage site known for its extensive colonial dry-stone walls and significant Indigenous heritage. We discover that these self-supervised features enable accurate mapping of stone wall structures on various computer vision backbones, including ResNet, Wide ResNet, and Vision Transformers (ViTs), through limited supervised fine-tuning. The proposed self-supervised pre-training strategy, DINO-CV, can learn structural and visual features from different terrain data views (i.e., MHS and VAT), thereby substantially reducing the need for labeled data in supervised fine-tuning. These findings highlight the potential of self-supervised learning on DEM derivatives for large-scale, automated mapping of stone walls in complex, vegetated, and data-scarce heritage environments.
@@ 2.1 DEM Visualization of Archaeological Landscapes
These LiDAR-derived measurements can be processed into digital elevation models (DEMs), which represent the elevation of the bare-earth surface by filtering out vegetation and non-ground elements. DEMs can provide detailed topographic information that enables the detection of subtle landscape features, including ancient structures, earthworks, and other archaeological remains. As a result, DEMs have shown particularly suitable for mapping and analyzing large, complex, and vegetated archaeological landscapes (Štular et al., 2012). To further enhance the visibility of these features, a variety of DEM visualization techniques have been developed, such as hillshading (Yoeli, 1967), slope analysis (Doneus and Briese, 2006), and sky-view factor (Kokalj et al., 2011). These methods aid researchers in identifying and interpreting landscape features that are often not discernible in standard aerial or satellite imagery.
@@ 2.2 Automated Mapping of Dry-Stone Wall Structures
Recent advances in deep learning have enabled the automated detection and mapping of archaeological features from remote sensing data. Previous study by Trotter et al. (2022) employed a supervised Fully convolutional Neural Networ (FCN) (Long et al., 2015) based on the U-Net architecture (Ronneberger et al., 2015) to automate the mapping of dry-stone walls in Denmark. Their model was trained on a dataset of digital terrain models (DTMs) with a spatial resolution of 0.4m, using ground-truth annotations of dry-stone wall locations. While the study demonstrated the potential of deep learning for this task, its applicability is limited by the availability of high-resolution terrain data, which is widely accessible in Denmark but not available in other regions. Furthermore, the Danish DTM dataset is updated only every five years, and the performance of the supervised model is constrained by the limited quantity of labeled training data.
@@
These studies highlight the potential of deep learning for the automated mapping of archaeological stone wall structures. However, a common limitation across existing approaches is the dependence on large quantities of high-quality labeled data to achieve strong segmentation performance. The generation of such datasets often requires expert knowledge and significant manual effort, which can be resource-intensive. Moreover, the limited availability of high-resolution terrain data in many regions restricts the generalizability and scalability of these methods to diverse archaeological contexts.
@@ 2.3 Self-supervision with Knowledge Distillation
An initial study proposed by Hinton et al. (2015), as a method to transfer knowledge from a large ensemble of models (teachers) into a single, compact student model. This approach enables efficient deployment while retaining the performance benefits of model ensembles. Recent SSL methods based on knowledge distillation, such as DINO (Caron et al., 2021) and BYOL (Grill et al., 2020), have adapted this teacher–student framework to a self-supervised, label-free setting by using momentum encoders and self-distillation to match intermediate feature representations and output distributions between teacher and student networks. These methods do not require labeled data and are well-suited for learning semantic and structural features from visual inputs. Recent extensions, DINOv2 (Oquab et al., 2024) and DINOv3 (Siméoni et al., 2025), push this paradigm to much larger scales. DINOv2 stabilizes training at scale and introduces an automated pipeline for curating large and diverse image datasets. DINOv3 further scales the approach to 7B parameters and trains on even larger curated datasets, emphasizing scalability and generalization. While recent extensions such as DINOv2 and DINOv3 push self-distillation toward foundation-scale models through large curated datasets and massive ViTs, these directions are beyond the scope of this work. Instead, our focus is on cross-view self-supervision tailored to high-resolution DEM derivatives, addressing archaeological mapping challenges where labeled data is scarce.
In parallel of SSL with self-distillation, contrastive self-supervised learning methods such as SimCLR (Chen et al., 2020a, b), MoCo (He et al., 2020), and MoCov2 (Chen et al., 2020c) learn visual features by maximizing agreement between different augmented views of the same image (positive pairs), while minimizing agreement between views of different images (negative pairs). This is typically achieved by contrasting representations in a latent embedding space. Although contrastive learning has shown strong performance across a range of visual tasks, it often requires large batch sizes or memory banks to provide a sufficient number of negative examples for effective training (Caron et al., 2021).
@@
Differently, we propose a self-supervised cross-view pre-training framework based on knowledge distillation paradigm to pre-train vision backbones on unlabeled DEM derivatives (i.e., MHS and VAT), which can improve stone wall mapping capability in complex and vegetated archaeological landscapes.
@@ 3 Methodology
Figure 1: The aerial image (Row 1), Multi-directional Hillshade (MHS, Row 2), and Visualization for Archaeological Topography (VAT, Row 3) of dry-stone walls in the Budj Bim landscape. From left to right, columns show: (1–2) multiple parallel thin walls and an isolated thin wall; (3–4) thick, zigzag walls and a straight wall partially occluded by vegetation; (5–6) roadside straight walls with gate openings, and a heavily occluded wall segment.
3.1 DEM Visualization of Budj Bim Landscapes
High-resolution DEM visualization techniques are useful visualization tools for mapping cultural heritage features in complex terrains, where archaeological structures such as dry-stone walls are often obscured in conventional aerial imagery. To enhance the visibility of low-lying structures and support automated stone wall mapping on archaeological landscape, we employ two DEM-derived visualization techniques: Multi-directional Hillshade (MHS) and Visualization for Archaeological Topography (VAT).
@@ 3.1 DEM Visualization of Budj Bim Landscapes
While aerial imagery provides high-fidelity semantic information, such as textures of objects and surrounding vegetation, it is often limited by occlusions caused by dense vegetation, particularly in rural or forested areas. In contrast, MHS and VAT visualizations derived from DEM data are unaffected by such occlusions and can effectively reveal dry-stone wall structures that are otherwise indiscernible in aerial views, as illustrated in Fig. 1. In detail, columns 1 and 2 of Fig. 1 highlight multiple thin, parallel stone walls and a solitary linear segment, which are clearly discernible in both MHS and VAT. Columns 3 and 4 display zigzag and straight thick walls that are partially or fully concealed by vegetation in the aerial view but remain visible in the DEM visualizations. Columns 5 and 6 show roadside walls with gate openings and heavily occluded segments.
Together, MHS and VAT offer complementary topographic features that are well-suited for detecting the diverse forms and different kinds of dry-stone walls across the Budj Bim cultural landscape. Their integration enables robust feature extraction and significantly enhances automated mapping performance where visual occlusion and subtle relief variations pose challenges to traditional image-based methods.
@@ 3.2 Self-supervised Cross-view Pre-training
Figure 2: (a) The DINO-CV framework for self-supervised pre-training. At each training iteration $k$, the student network $g_{\theta_{s}}$ receives a randomly sampled input from the collection of DEM-derived views (i.e., MHS and VAT), while the teacher network $g_{\theta_{t}}$ receives a fixed view (MHS). (b) The Pseudo Siamese Network used for supervised fine-tuning. This architecture comprises two teacher networks, each pre-trained with DINO-CV on a distinct input view (MHS and VAT), enabling the extraction of complementary features for stone wall mapping.
@@
The self-supervised cross-view pre-training framework, DINO-CV (Fig.2(a)), is designed to learn invariant visual and structural features across multiple high-resolution DEM derivatives to enhance mapping performance of models in the subsequent supervised fine-tuning. DINO-CV builds on the knowledge distillation paradigm introduced by Caron et al. (2021), where a student network $g_{\theta_{s}}$, comprising a vision backbone $f_{\theta_{s}}$ and a projection head $h_{\theta_{s}}$, such that $g_{\theta_{s}}=h_{\theta_{s}}\circ f_{\theta_{s}}$, is trained to match the output distribution of a teacher network $g_{\theta_{t}}$ by minimizing the cross-entropy between their temperature-scaled softmax outputs:
@@
During training, the student parameters $\theta_{s}$ are updated using stochastic gradient descent (SGD) to minimize the distillation loss. The teacher network however is not directly optimized; instead, its parameters $\theta_{t}$ are updated at each iteration using an exponential moving average (EMA) of the student’s parameters (Tarvainen and Valpola, 2017):
@@
3.3 Pseudo Siamese Network
In supervised fine-tuning, we adopt a Pseudo Siamese Network architecture (Fig. 2(b)) to effectively leverage the complementary representations of DEM derivatives. This architecture consists of two parallel teacher networks pre-trained with DINO-CV, each trained on a different data view. A Pseudo Siamese Network refers to a two-stream architecture where the networks share the same architecture but have independent weights. In our case, the MHS and VAT streams are parameterized by their respective teacher networks, each pre-trained separately with DINO-CV, allowing the two streams to specialize in their respective data view.

### v3

# Mapping Hidden Heritage: Self-supervised Pre-training on High-Resolution LiDAR DEM Derivatives for Archaeological Stone Wall Detection

@@ Abstract
Historic dry-stone walls hold significant cultural and environmental importance, serving as historical markers and contributing to ecosystem preservation and wildfire management during dry seasons in Australia. However, many of these stone structures in remote or vegetated landscapes remain undocumented due to limited accessibility and the high cost of manual mapping. Deep learning–based segmentation offers a scalable approach for automated mapping of such features, but challenges remain: 1.the visual occlusion of low-lying dry-stone walls by dense vegetation and 2.the scarcity of labeled training data. This study presents DINO-CV, a self-supervised cross-view pre-training framework based on knowledge distillation, designed for accurate and data-efficient mapping of dry-stone walls using Digital Elevation Models (DEMs) derived from high-resolution airborne LiDAR. By learning invariant geometric and geomorphic features across DEM-derived views, (i.e., Multi-directional Hillshade and Visualization for Archaeological Topography), DINO-CV addresses the occlusion by vegetation and data scarcity challenges. Applied to the Budj Bim Cultural Landscape at Victoria, Australia, a UNESCO World Heritage site, the approach achieves a mean Intersection over Union (mIoU) of 68.6% on test areas and maintains 63.8% mIoU when fine-tuned with only 10% labeled data. These results demonstrate the potential of self-supervised learning on high-resolution DEM derivatives for large-scale, automated mapping of cultural heritage features in complex and vegetated environments. Beyond archaeology, this approach offers a scalable solution for environmental monitoring and heritage preservation across inaccessible or environmentally sensitive regions.
@@
self-supervised learning , LiDAR , digital elevation model , cultural heritage , archaeological mapping
@@
[…]
[inst2]organization=Gunditj Mirring Traditional Owners Corporation,addressline=248 Condah Estate Road, city=Breakaway Creek, postcode=3303, state=Victoria, country=Australia
@@ 1 Introduction
Traditional methods of mapping dry-stone walls, such as manual surveys or aerial photography, can be impractical and labor-intensive, especially in remote or rugged terrain. Recent advancements in computer vision and deep learning have enabled the development of automated, scalable mapping techniques that can efficiently detect dry-stone walls from high-resolution imagery and LiDAR data (Trotter et al., 2022; Suh and Ouimet, 2023; Wang et al., 2024). Supervised deep neural networks, such as ConvNets (CNNs) (He et al., 2016), vision transformers (ViTs) (Dosovitskiy et al., 2021), and U-Net architectures (Ronneberger et al., 2015), can be trained on annotated datasets to learn the visual and spatial features of dry-stone walls, including their shape, texture, and spatial patterns. These deep learning-based mapping approaches substantially reduce the time and cost of mapping and offer the potential for large-scale registration of undocumented structures. However, the performance of these deep supervised models heavily depends on the availability of large, labeled datasets for model training, which are often difficult and costly to obtain. Another critical challenge is the visual occlusion of low-lying stone wall structures by dense vegetation in wild landscapes, which makes it difficult for models to accurately detect and localize these features.
@@
Grounded in these hypotheses, this study poses the following research question: How can self-supervised learning on high-resolution DEM derivatives improve the mapping performance of low-lying dry-stone walls in complex, vegetated environments? To answer this question, we explore the potential of self-supervised learning on high-resolution DEM derivatives for automated mapping of dry-stone walls in complex, vegetated environments. Hence, we propose a self-supervised cross-view pre-training strategy, DINO-CV, based on knowledge distillation to learn visual and structural features from multiple DEM derivatives generated from high-resolution aerial LiDAR data. The mapping performance of our approach is evaluated on the Budj Bim cultural landscape in Victoria, Australia, a UNESCO World Heritage site known for its extensive colonial dry-stone walls and significant Indigenous heritage. We demonstrate that these self-supervised features enable accurate and data-efficient mapping of stone wall structures on various computer vision backbones, including ResNet, Wide ResNet, and Vision Transformers (ViTs), through limited supervised fine-tuning. The proposed self-supervised pre-training strategy, DINO-CV, can learn structural and visual features from different terrain data views (i.e., MHS and VAT), thereby substantially reducing the need for labeled data in supervised fine-tuning. These findings highlight the potential of self-supervised learning on DEM derivatives for large-scale, automated mapping of stone walls in complex, vegetated, and data-scarce heritage environments.
@@ 2.1 DEM Visualization of Archaeological Landscapes
These LiDAR-derived measurements can be processed into digital elevation models (DEMs), which represent the elevation of the bare-earth surface by filtering out vegetation and non-ground elements. DEMs can provide detailed topographic information that enables the detection of subtle landscape features, including ancient structures, earthworks, and other archaeological remains. As a result, DEMs have proven particularly suitable for mapping and analyzing large, complex, and vegetated archaeological landscapes (Štular et al., 2012). To further enhance the visibility of these features, a variety of DEM visualization techniques have been developed, such as hillshading (Yoeli, 1967), slope analysis (Doneus and Briese, 2006), and sky-view factor (Kokalj et al., 2011). These methods aid researchers in identifying and interpreting landscape features that are often not discernible in standard aerial or satellite imagery.
@@ 2.2 Automated Mapping of Dry-Stone Wall Structures
Recent advances in deep learning have enabled the automated detection and mapping of archaeological features from remote sensing data. Previous study by Trotter et al. (2022) employed a supervised Fully convolutional Neural Network (FCN) (Long et al., 2015) based on the U-Net architecture (Ronneberger et al., 2015) to automate the mapping of dry-stone walls in Denmark. Their model was trained on a dataset of digital terrain models (DTMs) with a spatial resolution of 0.4m, using ground-truth annotations of dry-stone wall locations. While the study demonstrated the potential of deep learning for this task, its applicability is limited by the availability of high-resolution terrain data, which is widely accessible in Denmark but not available in other regions. Furthermore, the Danish DTM dataset is updated only every five years, and the performance of the supervised model is constrained by the limited quantity of labeled training data.
@@
Beyond archaeological applications, related research on drainage ditch mapping demonstrates the broader potential of combining deep learning with airborne laser scanning (ALS) for detecting subtle geomorphic features in vegetated environments. Lidberg et al. (2023) developed a deep neural network trained on ALS data and manually digitized ditch channels across ten regions in Sweden, achieving an 86% detection accuracy. Their study highlights how deep learning models can effectively identify fine-scale hydrological and geomorphic structures that are otherwise difficult to observe due to canopy cover or terrain complexity. This approach parallels the objectives of our work on dry-stone wall mapping, emphasizing the utility of high-resolution topographic data and learning-based feature extraction for large-scale, automated landscape analysis.
These studies together highlight the potential of deep learning for the automated mapping of archaeological stone wall structures. However, a common limitation across existing approaches is the dependence on large quantities of high-quality labeled data to achieve strong segmentation performance. The generation of such datasets often requires expert knowledge and significant manual effort, which can be resource-intensive. Moreover, the limited availability of high-resolution terrain data in many regions restricts the generalizability and scalability of these methods to diverse archaeological contexts.
@@ 2.3 Self-supervision with Knowledge Distillation
Hinton et al. (2015) introduced knowledge distillation as a method to transfer knowledge from a large ensemble of models (teachers) into a single, compact student model. This approach enables efficient deployment while retaining the performance benefits of model ensembles. Recent SSL methods based on knowledge distillation, such as DINO (Caron et al., 2021) and BYOL (Grill et al., 2020), have adapted this teacher–student framework to a self-supervised, label-free setting by using momentum encoders and self-distillation to match intermediate feature representations and output distributions between teacher and student networks. These methods do not require labeled data and are well-suited for learning semantic and structural features from visual inputs. Recent extensions, DINOv2 (Oquab et al., 2024) and DINOv3 (Siméoni et al., 2025), push this paradigm to much larger scales. DINOv2 stabilizes training at scale and introduces an automated pipeline for curating large and diverse image datasets. DINOv3 further scales the approach to 7B parameters and trains on even larger curated datasets, emphasizing scalability and generalization. While recent extensions such as DINOv2 and DINOv3 push self-distillation toward foundation-scale models through large curated datasets and massive ViTs, these directions are beyond the scope of this work. Instead, our focus is on cross-view self-supervision tailored to high-resolution DEM derivatives, addressing archaeological mapping challenges where labeled data is scarce.
In parallel with self-distillation SSL approaches, contrastive self-supervised learning methods such as SimCLR (Chen et al., 2020a, b), MoCo (He et al., 2020), and MoCov2 (Chen et al., 2020c) learn visual features by maximizing agreement between different augmented views of the same image (positive pairs), while minimizing agreement between views of different images (negative pairs). This is typically achieved by contrasting representations in a latent embedding space. Although contrastive learning has shown strong performance across a range of visual tasks, it often requires large batch sizes or memory banks to provide a sufficient number of negative examples for effective training (Caron et al., 2021).
@@
In contrast, we propose a self-supervised cross-view pre-training framework based on knowledge distillation paradigm to pre-train vision backbones on unlabeled DEM derivatives (i.e., MHS and VAT), which can improve stone wall mapping capability in complex and vegetated archaeological landscapes.
@@ 3 Methodology
Figure 1: Overview of the airborne LiDAR dataset covering 301 km2 of the Budj Bim Cultural Landscape in southwestern Victoria, Australia. The survey captured over 33 billion points using multi-return laser scanning, providing the high-resolution topographic data required for generating digital elevation models (DEMs) and derivative visualizations used in this study. The location of BudjbimArea (25 km2) dataset for fine-tuning is shown in a blue dashed bounding box, and the location of an unlabeled region used for qualitative demonstration is shown in a green dashed bounding box.
Figure 2: The aerial image (Aerial, Row 1), Multi-directional Hillshade (MHS, Row 2), and Visualization for Archaeological Topography (VAT, Row 3) of dry-stone walls in the Budj Bim landscape. From left to right, columns show: (1-2) multiple parallel thin walls and an isolated thin wall; (3-4) thick, zigzag walls and a straight wall partially occluded by vegetation; (5-6) roadside straight walls with gate openings, and a heavily occluded wall segment.
3.1 Airborne LiDAR Acquisition & Processing
To enable reliable detection of subtle archaeological features across the Budj Bim landscape, careful consideration was given to the timing and conditions of airborne LiDAR acquisition. The LiDAR survey was conducted in 2020 by the Department of Environment, Land, Water and Planning (DELWP), Victoria, on behalf of the Gunditjmara Traditional Owners, during the dry season and following bushfire events, when vegetation cover was minimal. These conditions maximized ground visibility and enhanced the detectability of subtle surface features. Capturing fine-scale archaeological traces such as dry-stone walls required high point densities, achieved through multiple overlapping flight passes at varying scan angles to ensure uniform surface coverage across the diverse terrain (Fig. 1).
Following data acquisition, the raw LiDAR point cloud was processed to isolate ground points from vegetation and other non-terrain elements. The Cloth Simulation Filtering (CSF) algorithm (Zhang et al., 2016) was adopted for this purpose due to its robustness in modelling complex ground surfaces under variable vegetation conditions. The CSF algorithm simulates a virtual cloth draped over an inverted point clouds, effectively distinguishing ground points by conforming to the underlying terrain surface while excluding elevated canopy structures.
Ground-classified points were subsequently interpolated to generate high-resolution DEMs at 10 cm spatial resolution. Rasterization and gridding were applied with carefully selected interpolation parameters to ensure smooth and accurate terrain representation. The resulting raster DEMs preserve the geomorphic characteristics of the landscape, including low-relief archaeological features that are often imperceptible in optical imagery. These processed DEMs served as the foundational layers for generating the DEM derivatives and visualizations used in the DINO-CV pre-training framework.
3.2 DEM Visualization of the Budj Bim Landscape
High-resolution DEM visualization techniques are useful visualization tools for mapping cultural heritage features in complex terrains, where archaeological structures such as dry-stone walls are often obscured in conventional aerial imagery. To enhance the visibility of low-lying structures and support automated stone-wall mapping on archaeological landscapes, the high-resolution DEMs are processed using Relief Visualization Toolbox (RVT) (Kokalj et al., 2011; Kokalj and Somrak, 2019) to generate two DEM-derived visualization methods: Multi-directional Hillshade (MHS) and Visualization for Archaeological Topography (VAT).
@@ 3.2 DEM Visualization of the Budj Bim Landscape
While aerial imagery provides high-fidelity semantic information, such as textures of objects and surrounding vegetation, it is often limited by occlusions caused by dense vegetation, particularly in rural or forested areas. In contrast, MHS and VAT visualizations derived from DEMs are unaffected by such occlusions and can effectively reveal dry-stone wall structures that are otherwise indiscernible in aerial views, as illustrated in Fig. 2. In detail, columns 1 and 2 of Fig. 2 highlight multiple thin, parallel stone walls and a solitary linear segment, which are clearly discernible in both MHS and VAT. Columns 3 and 4 display zigzag and straight thick walls that are partially or fully concealed by vegetation in the aerial view but remain visible in the DEM visualizations. Columns 5 and 6 show roadside walls with gate openings and heavily occluded segments.
Together, MHS and VAT offer complementary topographic features that are well-suited for detecting the diverse forms of dry-stone walls across the Budj Bim cultural landscape. Their integration enables robust feature extraction and significantly enhances automated mapping performance where visual occlusion and subtle relief variations pose challenges to traditional image-based methods.
@@ 3.3 Self-supervised Cross-view Pre-training
Figure 3: Overview of the DINO-CV architecture. (a) Self-supervised Cross-view Pre-training: The framework employs a student-teacher distillation setup. At each iteration $k$, the student network $g_{\theta_{s}}$ receives a randomly sampled input view $x^{(ik)}$ (e.g., VAT) from the collection $\mathcal{V}$, while the teacher network $g_{\theta_{t}}$ receives a fixed view $x^{t}$ (e.g., MHS). The student is optimized via stochastic gradient descent to match the teacher’s output, while the teacher parameters are updated via an Exponential Moving Average (ema) of the student weights. (b) Pseudo-Siamese Fine-tuning: For the downstream task, we instantiate a Pseudo-Siamese two-stream network initialized with weights from two separate teacher networks from DINO-CV pre-training (one specialized on MHS, one on VAT). These parallel encoders extract complementary geometric features which are fused via element-wise summation before being passed to the decoder for stone wall segmentation.
@@
The self-supervised cross-view pre-training framework, DINO-CV (Fig. 3a), is designed to learn invariant visual and structural features across multiple high-resolution DEM derivatives to enhance mapping performance of models in the subsequent supervised fine-tuning. DINO-CV builds on the knowledge distillation paradigm introduced by Caron et al. (2021), where a student network $g_{\theta_{s}}$, comprising a vision backbone $f_{\theta_{s}}$ and a projection head $h_{\theta_{s}}$, such that $g_{\theta_{s}}=h_{\theta_{s}}\circ f_{\theta_{s}}$, is trained to match the output distribution of a teacher network $g_{\theta_{t}}$ by minimizing the cross-entropy between their temperature-scaled softmax outputs:
@@
During training, the student parameters $\theta_{s}$ are updated using stochastic gradient descent (SGD) to minimize the distillation loss. The teacher network, however, is not directly optimized; instead, its parameters $\theta_{t}$ are updated at each iteration using an exponential moving average (EMA) of the student’s parameters (Tarvainen and Valpola, 2017):
@@
3.4 Pseudo-Siamese Network
In supervised fine-tuning, we adopt a Pseudo-Siamese Network architecture (Fig. 3b) to effectively leverage the complementary representations of DEM derivatives. This architecture consists of two parallel teacher networks pre-trained with DINO-CV, each trained on a different data view. A Pseudo-Siamese Network refers to a two-stream architecture where the networks share the same architecture but have independent weights. In our case, the MHS and VAT streams are parameterized by their respective teacher networks, each pre-trained separately with DINO-CV, allowing the two streams to specialize in their respective data view.